fond the period and amplitude of:
y=sin3x
y=-3sin2x
y=-2sin1/4x
y=5cos pi x/2

Answer by stanbon(75887)

fond the period and amplitude of:
y=sin3x
amp=3 ;period = 2(pi)/3 or 120 degrees
-------------------
y=-3sin2x
amp = 3 ; period = (2pi)/2 = pi or 180 degrees
------------------
y=-2sin(1/4)x
amp = 2 ; period = (2pi)/(1/4) = 8pi or 8*180 degrees
-------------------
y=5cos pi x/2
amp = 5 ; period = (2pi)/(pi/2) = 4 radians or 4(180/pi) degrees
---------------------
Chers,
Stan H.